1. Acetate buffer solution may be also prepared by the partial neutralization of HA with NaOH: 20 cm3 of the solution of NaOH, 0.1 mol dm-3 were mixed with 50 cm3 of the solution of HA with the same concentration, and water added up to 100 cm3. Calculate the equilibrium concentration of [A-] and pH of the final buffer solution.

2, Acetic acid (CH3COOH, abbrev. HA) is a very common weak acid with pKA = 4.75 (at 25oC and low ionic strength 0. 33) Calculate pH of the diluted solution of acetic acid with concentration c = 0.05 mol dm-3.
Answer #1

Q1:

Millimole of NaOH taken = 20ml × 0.1M = 2 mmol

Millimole of HA taken = 50ml × 0.1M = 5 mmol

HA + NaOH NaA + H2O

[A-] = 2mmol/100ml = 0.02M. (answer)

pH = pKa + log([A-]/[HA]) = 4.75 + log(2/100 ÷ 3/100)

= 4.75 + log(2/3)

= 4.5739

pH = 4.574 (answer)